Five resistors are connected in parallel with a resistance of 1 ohm, 2 ohms, 3 ohms, 4 ohms, and 5 ohms respectively. What is its equivalent resistance
1
Expert's answer
2021-10-18T11:05:55-0400

Solution;

For multiple resistors in parallel;

"\frac{1}{R_e}=\frac{1}{R_1}+\frac{1}{R_2}+\frac{1}{R_3}+..."

Substitute the give values;

"\frac{1}{R_e}=\frac11+\frac12+\frac13+\frac14+\frac15"

"\frac{1}{R_e}=\frac{137}{60}"

"R_e=\frac{60}{137}" ohms.